TMS Meeting Minutes, December 8, 2022

Attendees: Heejoo, John, Olga, Trenton, Yang (Zoom Meeting)

LBC Testing During Engineering Night:
  • 2022.12.02
  • Sx side had some instability when channel 21 was dropped
  • Intentionally dropped ch17, then ch20, then ch21, and then ch26; there were also other channels unintentionally dropped from std. dev. filtering
  • While ch17 was intentionally dropped, at the start of an exposure around 9:50 UT, there was a point where all channels dropped
  • During the period druing which ch21 was dropped, there was some anomalous behavior; random errors in alignment were encountered
  • When only ch26 was dropped, this random error went away, so it is clear that this is an issue specifically affecting ch21
  • Heejoo's next task will be to take a look at the distance values for ch21 and see how this channel contributes to the pose calculation (check the Jacobian sensitivity elements)
  • Note, ch27 wasn't used in the reference on this night, and it was consistently dropped throughout the night
  • Heejoo will dig into the Jacobian calculation to determine the root cause of the issue
Tonight
  • We may have a chance to run TMS around 3 AM MST or 2:30 AM tonight, if weather allows
  • Team meeting on polycom
  • Testing will go until the end of the night, at 6:15 AM at 12 degree twilight
  • This is all assuming that weather is good
  • Heejoo would like to conduct channel drop testing at lower elevation without filter changes
  • Heejoo will produce a written plan and share it with the team before the test, including which number should be dropped at which time
  • Olga will try to do better with how long she might take with MODS
  • If under pressure, would check out MODS in one hour; since no pressure, might take a little longer
  • May reconfigure for LBC around 2:15ish
  • There is forecast for strong gusts today, so may be unable to open the telescope (though today is better than yesterday); there are even three people on the roof cleaning ice off today
  • The team will test the passive mode switch on the python functionality tonight
  • Olga notes that VxWorks was restarted yesterday, and it is possible that the mirror was dropped hard; this means that many collimators may be misaligned
  • Heejoo would really like to install an accelerometer on the mirror to track when the collimator experience > 2 G loads (likely misalignment)
  • Yang suggested implementing a system with active alignment nobs for collimator adjustment following a bump
  • If there is an alignment issue, we should be able to see it tonight, even under closed-dome
Updating the Wiki Page:
  • Olga has recommended that we update the Wiki page with our most recent software version
  • The last engineering test, team ran the python version because of command line options
  • However, most users are currently using the perl version right now, so the webpage should be updated
  • Yang contends that the python version is working fine (except for one small problem with a shell printout warning that doesn't impact TMS functionality)
  • In the meantime, Yang suggests we allow people to continue using either the python version or the perl version
  • Yang to go over the wiki page next week to make edits and improve clarity
  • The reference setting program could use some command line options to change various parameters
  • Are we confident to allow people to use the python version? If so, Yang can write it into the wiki page and remove the original perl version, so as not to cause confusion
  • Things that people use most are the run TMS with one side only or to run TMS in passive mode (the python code does not currently have the passive switch, but Yang can add it; actually, it may already have the switch)
  • Yang will read over the wiki page to fix inconsistencies and bring things up to date; will also make sure the python verison supports everything that's needed right now
New LOFT Student and PhD to Join TMS:
  • Joon from the LOFT group will be joining the TMS effort and taking over for Trenton as he graduates
  • PhD student from INAF will also be joining TMS remotely to support
Expanding to include Tertiary Measurement:
  • Heejoo is talking with a fiber manufacturer about making a multicore fiber to run up to the telescope, replacing broken fibers and expanding past our current limit of 28 channels
  • Fiber price around $500, want two parts, so the total should come out to around $1000
  • Heejoo planning to write a proposal for funding to UofA and/or to LBTO to get this funded
  • During summer break, we may re-align or fix the broken fibers
  • We may also try the ASM on the Sx side

-- TrentonBrendel - 08 Dec 2022
Topic revision: r1 - 08 Dec 2022, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback